OpenStudy (anonymous):

WILL GIVE MEDAL AND BECOME FAN. Express in the form of " a+bi " \[\large \bf \sqrt{i}+\sqrt{-i}\]

OpenStudy (anonymous):

\[\large \bf \sqrt{i}+\sqrt{-i}\]

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

let me try let a+bi = sqrt i + sqrt(-i) square both sides

hartnn (hartnn):

a^2 - b^2 +2abi = i +i + 2sqrt (-i^2) (a^2-b^2 ) + (2ab)i = 2i + 2

hartnn (hartnn):

compare real and imaginary parts ab =1 a^2-b^2 = 2 something seems wrong...

hartnn (hartnn):

a^2 - b^2 +2abi = i -i + 2sqrt (-i^2) (a^2-b^2 ) + (2ab)i = 0 + 2

hartnn (hartnn):

yeah, so ab = 0 a^2-b^2 = 2 if b =0 , a^2 =2, s = sqrt 2

hartnn (hartnn):

hence sqrt i + sqrt (-i) = sqrt 2 + 0i :)

hartnn (hartnn):

wolf agrees :) http://www.wolframalpha.com/input/?i=sqrt+i+%2B+sqrt%7B-i%7D

hartnn (hartnn):

ask if any doubt in any step

OpenStudy (anonymous):

oh!! thank you .. @hartnn

OpenStudy (anonymous):

i got it